### v2

# Larger Is Not Always Better: Exploring Small Open-source Language Models in Logging Statement Generation

Abstract.
Developers use logging statements to create logs that document system behavior and aid in software maintenance. As such, high-quality logging is essential for effective maintenance; however, manual logging often leads to errors and inconsistency. Recent methods emphasize using large language models (LLMs) for automated logging statement generation, but these present privacy and resource issues, hindering their suitability for enterprise use. This paper presents the first large-scale empirical study evaluating small open-source language models (SOLMs) for automated logging statement generation. We evaluate four prominent SOLMs using various prompt strategies and parameter-efficient fine-tuning techniques, such as Low-Rank Adaptation (LoRA) and Retrieval-Augmented Generation (RAG). Our results show that fine-tuned SOLMs with LoRA and RAG prompts, particularly Qwen2.5-coder-14B, outperform existing tools and LLM baselines in predicting logging locations and generating high-quality statements, with robust generalization across diverse repositories. These findings highlight SOLMs as a privacy-preserving, efficient alternative for automated logging.
Software Logging, Logging Statement, Logging Text, Logging Practice, Large Language Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Maintaining Software
1. Introduction
Logs are textual records generated during software execution to capture runtime events, states, and contextual information (Yu et al., 2023b). A typical log statement consists of three components: a verbosity level, logging variables, and logging texts (He et al., 2021; Jiang et al., 2024b). In particular, as the example shown below, the logging level (e.g., debug) reflects the event’s severity; the logging variables (e.g., terminalState) hold critical run-time data about system states; meanwhile, the logging text (e.g., Stopping the checkpoint services with state) offers a static explanation of the system’s actions.
log.debug("Stopping the checkpoint services with state { }.", terminalState);
High-quality logs provide actionable insights for developers to diagnose failures, optimize system behavior, and ensure reliability. However, the absence or inadequacy of logging statements can severely hinder downstream tasks such as anomaly detection (Zhang et al., 2022; Li et al., 2020b; Liu et al., 2023) and failure diagnosis (Yu et al., 2023a; Jiang et al., 2025; Huang et al., 2024b), leading to prolonged debugging cycles and increased maintenance costs. Consequently, the strategic placement and content of log statements directly influence the effectiveness of software maintenance and evolution (Soldani and Brogi, 2022; Huang et al., 2024a).
Despite their critical role in software maintenance, producing high-quality logs manually is far from straightforward for developers. First, the absence of universal logging guidelines leads to inconsistent practices, where log quality, granularity, and utility vary widely based on individual developers’ expertise. This inconsistency complicates log analysis and reduces their diagnostic value (Chen and Jiang, 2019). Second, developers face the difficult task of balancing log quantity and quality: over-logging burdens systems with excessive data, while under-logging risks missing critical information (Yao et al., 2018). Third, the cognitive and time-intensive nature of manual logging further exacerbates these issues, as developers must anticipate complex system behaviors and failure points, often resulting in logs that are either too vague or overly detailed (Hassani et al., 2018). Additionally, maintaining log relevance over time is challenging, as software evolution can render existing logs obsolete (Zhong et al., 2025). These challenges highlight the need for automated logging statement generation (hereafter referred to simply as ‘automated logging’) solutions that can consistently produce high-quality logs.
To address these challenges, at the early stage, researchers have explored automated logging techniques by decomposing the problem into sub-tasks, such as where-to-log (identifying code locations for logging) (Zhao et al., 2017; Li et al., 2020a), what-to-log (generating static content and dynamic variables) (Ding et al., 2022; Liu et al., 2019), and log-level suggestion (Heng et al., 2024; Li et al., 2021a; Liu et al., 2022). However, these fragmented approaches lack integration into a unified, end-to-end framework for generating complete logging statements. Recent advances in pre-trained language models (LMs) have opened new avenues for automated logging. LANCE (Mastropaolo et al., 2022) and LEONID (Mastropaolo et al., 2024) pioneered the use of sequence-to-sequence models (e.g., T5 (Raffel et al., 2020)) to generate logging statements directly from code contexts. Subsequent tools, such as Fastlog (Xie et al., 2024), Unilog (Xu et al., 2024), and SCLogger (Li et al., 2024b), further leveraged the large language model (LLM) to improve logging quality. In particular, Unilog and SCLogger adopted prompt-based methods with LLMs such as GPT-3.5 and Codex, achieving state-of-the-art performance by harnessing the code comprehension and natural language generation capabilities of LLMs.
Despite their effectiveness, LLM-based logging tools (Xie et al., 2024; Xu et al., 2024; Li et al., 2024b) face several limitations in enterprise settings.
  • Privacy Risks. Sending proprietary code to commercial LLM APIs, such as OpenAI’s, risks exposing sensitive intellectual property (Yao et al., 2024). For instance, Samsung banned employee use of ChatGPT and other generative AI tools after an engineer accidentally leaked sensitive source code to ChatGPT (Kharpal, 2023).
  • Style Misalignment. LLMs, trained on generic datasets, struggle to generate logs that align with enterprise-specific logging styles, such as unique verbosity levels or error prioritization requirements, limiting their utility for organizational needs (He et al., 2022; Gu et al., 2022).
To address these issues, enterprises often consider fine-tuning and deploying open-source LLMs in private environments. However, this approach demands substantial computational resources, including thousands of GPU hours, which is impractical for many resource-constrained organizations (Chien et al., 2023; Imani et al., 2024). These limitations necessitate more accessible and efficient solutions for automated logging.
Small open-source language models (SOLMs), defined as open-source models with fewer than 14 billion parameters, have gained traction in software engineering tasks, such as program repair (Silva et al., 2023) and comment rectification (Rong et al., 2024), making them a promising solution for automated logging. By deploying SOLMs locally on consumer-grade hardware, such as an A100 GPU, enterprises can safeguard proprietary code, eliminating privacy risks associated with commercial APIs (Yang et al., 2024b). Moreover, SOLMs require significantly fewer computational resources, reducing costs and aligning with sustainable computing goals (Imani et al., 2024). Additionally, SOLMs can be fine-tuned on enterprise-specific codebases to produce logs that meet unique organizational requirements, such as specific formats or compliance standards (Lu et al., 2023; Rong et al., 2024). For resource-constrained enterprises, SOLMs offer a practical balance of privacy, efficiency, and adaptability, making their application to automated logging highly attractive. However, to the best of our knowledge, no studies have systematically explored the effectiveness of SOLMs in automated logging.
To fill this significant gap, in this paper, we conduct an empirical study on four prominent SOLMs, namely LLaMA (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), CodeLLaMA (Roziere et al., 2023), and Qwen2.5coder (Hui et al., 2024), to explore the potential utility of SOLMs in automatic generation of logging statements. We pose four research questions to comprehensively assess the potential of SOLMs.
RQ1: What are the most effective prompting strategies for using SOLMs in logging generation? Different prompting techniques (e.g., in-context learning (ICL) (Brown et al., 2020), chain-of-thought (COT) (Wei et al., 2022), retrieval-augmented generation (RAG) (Lewis et al., 2020)) can influence the performance of SOLMs without the need for retraining. Gaining insight into their effects can help improve the generation of logging statements across different contexts.
Result. RAG outperforms ICL and COT, significantly enhancing the performance of logging automation task.
RQ2: What is the best tuning strategy using SOLMs for automated logging? Many strategies such as parameter-efficient fine-tuning (PEFT) techniques (Houlsby et al., 2019; Hu et al., 2022), model size, and model type may influence the efficacy. Thus, we further investigate the extent of their impact, which may offer insights into the optimal selection of strategies for enhancing SOLM performance.
Result. LoRA (Hu et al., 2022) demonstrates the most consistent and superior results when fine-tuning with PEFT techniques. For models with more than 3B parameters, performance in generating logging statements improves with more parameters, but the increased computational costs indicate a trade-off between performance and resource costs. The instruct variant of the SOLM model outperforms its base counterpart, benefiting from its instruction-tuned foundation.
RQ3: How effectively do SOLMs compare to existing methods and LLM baselines in automated logging? Upon recognizing the optimal strategies for employing SOLMs, we aim to investigate the performance of SOLMs in automated logging compared to existing methods and the prominent LLMs.
Result. Fine-tuned SOLMs, particularly Qwen2.5-coder-14B, outperform both existing methods and LLMs across all evaluated metrics, demonstrating superior logging location accuracy and statement quality. The result of the LLM-based judger further supports the high quality of SOLM-generated statements.
RQ4: Can SOLMs generalize logging statement generation across diverse code repositories? Logging practices vary across projects. Evaluating SOLM generalization ability on diverse repositories ensures their applicability in real-world development environment.
Result. SOLMs demonstrate strong generalization capabilities in automated logging, maintaining high performance across diverse, unseen repositories. Similar logging practices, such as those shared among Apache open-source projects, significantly improve the cross-project generalization ability of SOLMs.
Contributions. To sum up, in this paper, we make the following contributions:
  • We conduct the first large-scale empirical study assessing the effectiveness of SOLMs for automated logging. Our findings demonstrate their capability to produce contextually accurate and syntactically correct logging statements that rival or exceed the performance of existing specialized methods and LLMs.
  • We investigate and identify effective strategies for optimizing SOLM performance in the context of automated logging. This includes demonstrating that RAG significantly enhances performance and that Low-Rank Adaptation stands out as a highly effective PEFT technique.
  • We showcase the practical advantages of fine-tuned SOLMs, including their robust generalization capabilities across diverse and previously unseen code repositories. This research highlights that SOLMs can maintain high performance in varied settings and offer benefits such as local deployment for data privacy and alignment with enterprise-specific logging styles.
  • To facilitate further research and encourage practical adoption in the field of automated logging, we publicly release our source code, datasets, and comprehensive experimental results (Renyi Zhong, Yichen Li, Guangba Yu, Wenwei Gu, Jinxi Kuang, Yintong Huo, Michael R. Lyu, 2025).
Paper Orgenizations. Section 2 discusses the background. Section 3 describes the experimental design of our study. Section 4 presents the experimental results. Section 5 introduces the advantages of using SOLMs for automated logging and potential future work directions. Section 6 discusses threats to validity. Section 7 introduces the related work. Section 8 concludes the paper.
2. Background
2.1. Problem Definition
Figure 1. Task formulation: given a method which missing a logging statement, the model is asked to automated generate a logging statement.
This paper focuses on the automated logging task (i.e., where-to-log + what-to-log), which to some extent can be viewed as a code editing problem: when presented with lines of code, usually corresponding to a method, the generator’s task is to identify both the precise location for logging, referred to as the logging point, and the complete logging statement (i.e., level, variables, and text). The predicted logging point should match the one that was originally present in the source file before being removed, and the predicted logging statement itself should closely resemble the excised original. Figure 1 provides a visual example of this task, showing how a proficient logging statement generator would intelligently incorporate LOG.info("Generating " + totalRows + " using " + numSplits); at line 4. It is important to highlight that this task is distinctly separate from the comprehensive empirical investigation conducted by Li et al. (Li et al., 2024a), which predominantly examines the question of what-to-log but lacks the consideration of where-to-log.
2.2. Large Language Models
The evolution of language models in recent times can be divided into three transformative phases. Initially, there were neural language models (NLMs), followed by the phase of pre-trained language models (PLMs), and the current era sees the prominence of LLMs. Pre-trained models such as CodeT5 (Wang et al., [n. d.]) and PLBART (Ahmad et al., 2021) have achieved noteworthy success in software engineering applications primarily due to task-specific pre-training processes. However, large language models have brought about a revolutionary change in the field due to their immense parameter counts, often exceeding 10 billion, and their comprehensive pre-training data. These models, unlike their pre-training predecessors, exhibit emergent capabilities that allow them to achieve robust performance across a wide range of tasks without necessitating pre-training tailored to specific tasks (Geng et al., 2024). This quality substantially diminishes the requirement for resource-heavy training sessions. Within the realm of software engineering, large language models are mainly categorized into two groups. Unified large language models, such as GPT-4o and LLaMA, which are designed to integrate natural language and code corpora, whereas code-specific large language models, like StarCoder (Li et al., 2023a) and CodeLlama (Roziere et al., 2023), are developed for specialization in tasks centered around coding.
Current methodologies for leveraging LLMs fall into two paradigms: prompt-based and tuning-based. Prompt-based methods exploit the zero-shot or few-shot capabilities of massive LLMs (e.g., GPT-4) through carefully engineered prompts, avoiding the need for explicit training data. They use techniques such as in-context learning (ICL) and chain-of-thought (COT). Conversely, fine-tuning approaches focus on parameter-efficient adaptation (e.g., prompt tuning, prefix tuning, and low-rank adaptation) to tailor smaller-scale LLMs for domain-specific tasks. These techniques freeze base model parameters while training minimal additional components, achieving performance comparable to full-parameter tuning at reduced computational costs.
2.3. LLM Applications in Software Engineering Task
Researchers have conducted in-depth investigations into the use of Large Language Models (LLMs) in a variety of software engineering tasks, including but not limited to code completion (Izadi et al., 2022, 2024), vulnerability detection (Sun et al., 2024; Zhao et al., 2024), program repair (Wei et al., 2023; Lin et al., 2024), and test generation (Chen et al., 2024; Liu et al., 2024a). These studies highlight the versatility and adaptability of LLMs in effectively tackling a range of software engineering challenges.
In certain tasks, approaches that utilize prompts have been shown to yield superior outcomes (Imani et al., 2024; Jiang et al., 2024a). Pan et al., for instance, studied the efficiency of LLMs in the context of code translation (Pan et al., 2024). Within the range of models assessed, which encompassed both SOLMs and GPT-4, the top-performing SOLM, known as StarCoder, managed to reach a successful translation rate of 14.5%, in contrast to the success rate of 47.3% achieved by GPT-4. In contrast, SOLMs have been shown to achieve comparable outcomes in specific domains. For instance, Tian et al. (Tian et al., 2024) observed an F1-score of 86.58% using UniXCoder for the task of detecting equivalent mutants, providing a notable contrast to the performance of GPT-4, which recorded an F1-score of approximately 55.90%. Furthermore, when employing Vicuna 13B in conjunction with the innovative LogPrompt strategy, the performance was found to be on par with that of GPT-4, as reported in (Liu et al., 2024b).
3. Experimental Design
Figure 2 illustrates the overview of our study design. Initially, based on the AL-Bench dataset (Tan et al., [n. d.]), we construct the fine-tuning, valid and test datasets. Then we investigate SOLMs for automated logging through four research questions (RQs). RQ1 investegates the most effective prompt strategies for using SOLMs in this task. RQ2 aims to determine the optimal fine-tuning strategies by PEFT techniques, model sizes, and model types. Following that, RQ3 compares the performance of fine-tuned SOLMs against existing methods and LLM baselines. Finally, RQ4 assesses the ability of SOLMs to generalize their logging statement generation capabilities across diverse unseen code repositories.
Figure 2. The overview of our experimental design with four research questions.
3.1. Dataset Preparation
Table 1. Statistics of source repositories of AL-Bench (Tan et al., [n. d.]).
Index Dataset Domains #Stars #Forks #LOLS
R1 Dbeaver Database Management 43.3k 3.7k 1.7k
R2 Dolphinscheduler Task Scheduling 13.5k 4.8k 1.9k
R3 Doris High Performance Database 13.6k 3.4k 2.9k
R4 Flink Data Processing 24.8k 13.5k 3.2k
R5 Hadoop Distributed Storage 15.1k 9.0k 16.0k
R6 Kafka Messaging Systems 30.0k 14.3k 3.4k
R7 Keycloak Identity and Access Management 26.9k 7.2k 1.0k
R8 Pulsar Messaging Systems 14.6k 3.6k 7.2k
R9 Thingsboard IoT Platform 18.7k 5.5k 2.7k
R10 Zookeeper Distributed Coordination 12.5k 7.3k 1.9k
3.1.1. Studied dataset.
To evaluate the performance of automated logging, we select the most recently proposed AL-Bench (Tan et al., [n. d.]). This dataset, containing 42,224 logging statements from 10 widely used projects, was specifically constructed using stringent criteria (e.g., ¿=10k stars, ¿=1k logging statements, ¿=500 log-related issues per project) to ensure the inclusion of well-maintained repositories where developers prioritize high-quality logging statements. Furthermore, AL-Bench intentionally incorporates projects from diverse domains, such as database management, task scheduling, distributed systems, messaging, and IoT platforms, to ensure comprehensive coverage of various real-world logging requirements and practices, making it a robust benchmark. Measures like using recent project versions and standardized code formatting were also employed during its creation to mitigate potential data contamination risks from pre-training corpora. Table 1 provides further statistical details on these source repositories.
3.1.2. Pre-processing and dataset construction.
To align with the research objectives of this study and accommodate the requirements of certain baseline models, we performed several adjustments to the original dataset. Firstly, recognizing that some baseline models have an input token limit, we filtered out data instances where the input code exceeded 512 tokens. Secondly, the original construction of AL-Bench could generate multiple data points from a single Java function if it contained multiple logging statements, with each data point representing one specific automated logging case. To prevent potential data leakage, where highly similar code snippets from the same source file might inadvertently appear across different dataset splits (e.g., fine-tuning and testing), we implemented a file-level splitting strategy. This approach ensures that all data instances originating from the same Java file are strictly allocated to only one of the fine-tuning, validation, or test sets. After applying these pre-processing steps, we obtained a final dataset comprising 33,224 instances. We then partitioned this dataset into fine-tuning, validation, and test sets, targeting an 8:1:1 ratio. Because our file-level splitting strategy required keeping all instances from a single file within the same set, the resulting distribution was approximate. The final fine-tuning set contains 26,713 instances, the validation set contains 3,508 instances, and the test set contains 3,003 instances.
3.2. Studied Models
In our study, we investigate the performance of the following SOLMs for logging statement generation. These models have been widely adopted in the literature related to SE tasks, including:
  • LLaMA3 (Grattafiori et al., 2024) is Meta’s latest LLM and refines the LLaMA 2 framework. It stands as a prominent open source LLM used in numerous software applications. Trained on an extensive and varied dataset far surpassing its predecessor, LLaMA 3 exhibits significantly improved proficiency in reasoning, code generation, and instruction adherence.
  • Mistral (Jiang et al., 2023) is noted for its efficiency and performance, utilizing Grouped-Query Attention (GQA) and Sliding Window Attention (SWA) for faster inference and broader context handling, and demonstrates robust general abilities and notable coding skills.
  • CodeLlama (Roziere et al., 2023) is a series of LLMs specialized in generating and completing code, based on the LLaMA2 framework. These models are initially trained using a dataset of 500 billion code tokens and subsequently refined to manage extended context effectively.
  • Qwen2.5-coder (Hui et al., 2024) is a code-specialized version of the Qwen2.5 (Yang et al., 2024a) family, which inherits Qwen’s multilingual capabilities and architectural improvement. While demonstrating strong and comprehensive coding abilities, it also possesses good general and mathematical skills.
3.3. Baselines
To evaluate SOLMs performance, we select the baselines by conducting a literature review of relevant papers published in SE venues. From this, we find the following methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2024), UniLog (Xu et al., 2024), and FastLog (Xie et al., 2024). Additionally, we examine several LLM baselines, including general-purpose LLMs (Claude3.7-sonnet, GPT4o, LLaMA3.1-405b), and a code-specific LLM (Deepseek-coder-v3). In relation to the methodologies for prompting, we derive our approach from the study (Gao et al., 2024) and incorporate four different strategies: instruction prompting (base), in-context learning (ICL), retrieval-augmented generation (RAG) and chain of thought (CoT). The instruction prompting strategy involves directly prompting LLMs to generate logging statements using identical inputs as those provided to SOLMs, without any supplementary data. The ICL approach consists of providing one random example before the main query to assist the model in grasping the nature of the task more effectively. In the RAG approach, rather than selecting example randomly, we retrieve precise examples from the valid dataset to accommodate various input samples. We specifically employ the BM25 algorithm to identify the case within the valid dataset that most closely resembles the query sample in terms of similarity. The details of the prompts are shown in Figure 3.
3.4. Strategies for Parameter-Efficient Fine-Tuning
We examine how the following PEFT strategies influence the performance of SOLMs when automated logging.
Prefix tuning (Li and Liang, 2021), is a PEFT strategy designed to adapt LLMs to specific downstream tasks while keeping the original model parameters entirely frozen. Instead of modifying the model’s weights, it introduces a small set of trainable continuous vectors, known as the ”prefix”, which are prepended to the key and value sequences within the multi-head attention mechanisms of the Transformer architecture, typically applied to the topmost $L$ layers. Specifically, for a given layer $l$, trainable prefix matrices $\mathbf{P}_{k}\in\mathbb{R}^{K\times C}$ and $\mathbf{P}_{v}\in\mathbb{R}^{K\times C}$ (where $K$ is the prefix length, a key hyperparameter, and $C$ is the hidden dimension) are concatenated with the original key ($\mathbf{K}_{l}\in\mathbb{R}^{M\times C}$) and value ($\mathbf{V}_{l}\in\mathbb{R}^{M\times C}$) matrices derived from the $M$ input tokens, forming augmented matrices $\mathbf{K}^{\prime}_{l}=[\mathbf{P}_{k};\mathbf{K}_{l}]$ and $\mathbf{V}^{\prime}_{l}=[\mathbf{P}_{v};\mathbf{V}_{l}]$. During fine-tuning, only the parameters comprising these prefix matrices ($\mathbf{P}_{k},\mathbf{P}_{v}$ across the selected layers) are optimized via gradient descent, learning task-specific representations that effectively steer the frozen model’s attention and subsequent computations towards generating appropriate outputs for the target task. This approach significantly reduces the number of trainable parameters compared to full fine-tuning, requires storing only the small prefix parameters per task, and avoids catastrophic forgetting by leaving the core model untouched.
Prompt tuning (Lester et al., 2021) offers an even more lightweight approach by confining trainable parameters exclusively to continuous prompt embeddings added only at the input layer, while freezing the entire pre-trained model, including its word embedding table. This method prepends a sequence of $K$ learnable prompt embeddings, represented by a single trainable matrix $\mathbf{P}_{\text{emb}}\in\mathbb{R}^{K\times C}$ (where $K$ is the prompt length and $C$ is the model’s embedding dimension), to the original sequence of $M$ input token embeddings $\mathbf{E}\in\mathbb{R}^{M\times C}$, yielding an augmented input sequence $\mathbf{E}^{\prime}=[\mathbf{P}_{\text{emb}};\mathbf{E}]$. This combined sequence $\mathbf{E}^{\prime}$ is then fed directly into the first layer of the frozen Transformer backbone. During the fine-tuning process, only the parameters of the prompt embedding matrix $\mathbf{P}_{\text{emb}}$ are updated. The core idea is that these learned continuous vectors act as a ”soft prompt” or task instruction, conditioning the frozen model’s behavior without requiring any internal modifications. Prompt Tuning demonstrates significant efficiency regarding parameter usage, typically necessitating the update of fewer than 0.1% of the total model parameters. This characteristic renders it highly efficient in terms of both storage and computation, especially in multi-task contexts.
Lora (Hu et al., 2022) provides a distinct PEFT mechanism based on the hypothesis that the change in weights during model adaptation has a low intrinsic rank. Instead of adding prefix or prompt tokens, LoRA freezes the original pre-trained weights $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ of selected layers (commonly the query, key, value, and output projection matrices in self-attention, and sometimes feed-forward layers) and injects trainable, rank-decomposition matrices in parallel. Specifically, the weight update $\Delta\mathbf{W}$ is approximated by the product of two smaller, low-rank matrices: $\mathbf{W}_{\text{down}}\in\mathbb{R}^{d\times r}$ and $\mathbf{W}_{\text{up}}\in\mathbb{R}^{r\times k}$, where the rank $r$ is a crucial hyperparameter significantly smaller than the original dimensions ($r\ll\min(d,k)$). The modified forward pass for an input $\mathbf{x}$ computes the output as the sum of the original path and the adapter path: $\mathbf{h}_{\text{adapted}}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\alpha(\mathbf{W}_{\text{down}}\mathbf{W}_{\text{up}})\mathbf{x}$, where $\alpha$ is often a constant scaling factor (like $\alpha/r$). During fine-tuning, only the parameters of $\mathbf{W}_{\text{down}}$ (typically initialized randomly) and $\mathbf{W}_{\text{up}}$ (typically initialized to zero) are optimized. A significant advantage of LoRA is that the learned adapter weights can be mathematically merged with the original weights ($\mathbf{W}=\mathbf{W}_{0}+\mathbf{W}_{\text{down}}\mathbf{W}_{\text{up}}$) after training, resulting in a single weight matrix per adapted layer and incurring zero additional inference latency compared to the original model, while still offering substantial parameter savings during training and allowing easy task switching by loading different adapter pairs.
QLora (Dettmers et al., 2023) represents a significant advancement in memory-efficient fine-tuning, specifically designed to make the adaptation of extremely large language models feasible on commodity hardware with limited VRAM. It ingeniously combines low-precision quantization of the base model with the LoRA technique. The core strategy involves loading the massive pre-trained base model $\mathbf{W}_{0}$ with its weights quantized to a very low bit-format, most notably 4-bit NormalFloat (NF4), a data type empirically shown to be effective for normally distributed weights, and keeping these quantized weights $Q(\mathbf{W}_{0})$ frozen. Standard LoRA adapters, consisting of low-rank matrices $\mathbf{W}_{\text{down}}$ and $\mathbf{W}_{\text{up}}$, are then introduced parallel to these quantized layers, but crucially, these adapter weights are maintained and trained in a higher precision format, typically BFloat16, to preserve adaptation capacity. The forward pass thus involves computations using the low-precision base model and the higher-precision adapters: $\mathbf{h}_{\text{adapted}}\approx Q(\mathbf{W}_{0})\mathbf{x}+\alpha(\mathbf{W}_{\text{down}}\mathbf{W}_{\text{up}})\mathbf{x}$. To further minimize memory bottlenecks, QLoRA incorporates innovations like double quantization and paged optimizers. By drastically reducing the memory footprint of the base model weights, activations (due to lower precision), and optimizer states, QLoRA enables fine-tuning models with tens or hundreds of billions of parameters on single consumer GPUs, while aiming to retain the task performance levels achieved by full-precision LoRA.
3.5. Evaluation method
3.5.1. Tradictional evaluating metrics
Considering earlier research (Li et al., 2024a; Tan et al., [n. d.]), we assess the performance of automated generation of logging statements by focusing on four aspects: the logging point, the logging levels, the logging text, and the logging variables. While each of these components highlights distinct aspects of system runtime information, they collectively serve as essential and complementary resources that aid engineers in analyzing and understanding system behaviour.
Logging point: We use position accuracy (PA) to evaluate the performance of logging location. To quantify PA, we compare the predicted locations of logging statements against their ground truth positions in the source code. This metric is formally defined as the ratio of correctly positioned logging statements ($LS_{position\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{position\_correct}}{LS_{all}}$.
Logging level: We use the Level Accuracy (LA) and Average Ordinal Distance Score (AOD) for evaluating the prediction of logging levels. Given the significant semantic differences between these levels and their implications for system monitoring and maintenance, we rigorously assess LA by comparing predicted log levels against their ground truth values in the source code. This metric is formally defined as the ratio of correctly predicted log levels ($LS_{level\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{level\_correct}}{LS_{all}}$. AOD evaluates how closely the current logging level aligns with the recommended logging level for each specific logging statement, as detailed in (Li et al., 2021a). The formula to calculate AOD is given by: $AOD=\frac{\sum_{i=1}^{N}(1-\frac{Dis(a_{i},s_{i})}{MaxDis(a_{i})})}{N}$, where $N$ represents the total number of logging statements in consideration. The term $MaxDis(a_{i})$ is used to denote the maximum potential distance for the actual log level $a_{i}$.
Static logging text: Same as previous study (Li et al., 2024a; Mastropaolo et al., 2022; Ding et al., 2022) , our evaluation of static logging texts is conducted through the application of two metrics commonly employed in the domain of machine translation: BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004). These metrics, grounded in n-gram analysis, are instrumental in assessing the degree of similarity between log messages that are generated computationally and those authored by developers. They provide a normalized score continuum from 0 to 1, with elevated scores indicative of a closer resemblance. In our methodology, we specifically implement various forms of these metrics, identified as BLEU-4 and ROUGE-L.
Dynamic logging variables: We use Precisely Match Rate (PMR) and F1 to evaluate dynamic logging variables. PMR ensures consistency in the capture of variable runtime data—a critical aspect of log effectiveness. We extract dynamic variables from both reference implementations and predicted logging statements, then perform exact matching to evaluate correspondence. PMR is formally defined as the ratio of exactly matched dynamic variables ($LS_{variable\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{variable\_correct}}{LS_{all}}$. Moreover, consider each logging statement and let us define $S_{ud}$ as the set of variables included in the generated logging statement, while $S_{gt}$ pertains to the set of variables present in the actual logging statement. In our analysis, we determine and present the following metrics: the precision, which is the ratio of variables from the updates that accurately match those in the actual logging ($precision=\frac{S_{ud}\cap S_{gt}}{S_{ud}}$); the recall, which indicates the fraction of actual variables correctly anticipated by the model ($recall=\frac{S_{ud}\cap S_{gt}}{S_{gt}}$); and, lastly, their harmonic mean, expressed as the F1 score ($F_{1}=2*\frac{Precision*Recall}{Precision+Recall}$(Li et al., 2024a).
3.5.2. LLM-as-a-judge
The evaluation of automatically generated logging statements, drawing upon our experimental findings and established prior work, conventionally proceeds by assessing distinct modules such as placement, verbosity level, and textual content. However, it is increasingly evident that certain prevalent metrics do not adequately capture the nuanced performance aspects of generated logging statements (Roy et al., 2021; Gros et al., 2020). For instance, in the assessment of static text components, metrics like BLEU-4 and ROUGE-L are confined to lexical similarity, largely overlooking crucial semantic congruity. This limitation presents a formidable challenge in establishing a unified and robust methodology for evaluating the overall quality of automatically generated logging statements.
Recently, research within the NLP domain has explored the application of LLM to appraise the quality of LLM-generated content, known as ”LLM-as-a-judge”. While human evaluation remains a reliable way, its inherent drawbacks of being time-consuming and cost-intensive run counter to the objectives of automated evaluation. Consequently, researchers are increasingly investigating methods to prompt or train LLMs to align with human evaluative preferences, thereby offering a scalable alternative to manual assessment. Supporting this direction, an empirical study by Wang et al. (Wang et al., 2025) has demonstrated the efficacy of the LLM-as-a-judge approach across various SE tasks. Their findings indicate that output-based evaluation methods, when coupled with state-of-the-art LLMs, yield optimal performance irrespective of the specific inference strategies employed. Informed by these advancements, this paper adopts the LLM-as-a-judge methodology to augment the quality assessment of automatically generated logging statements.
Specifically, we select three LLMs recognized for their superior performance in code-related tasks: Claude3.7-Sonnet, Deepseek-coder-v3, and GPT-4o. We choose this model due to their robust code comprehension and generation capabilities, which are critical for assessing the nuanced quality of logging statements. The evaluation process involves providing each LLM judge with the input code context, the generated logging statement from model, and the corresponding ground truth logging statement. The judges assign scores ranging from 0 to 3, where higher scores indicate greater alignment with the ground truth in terms of logging point accuracy, level appropriateness, static text quality, and dynamic variable correctness. For each generated logging statement, the final score is the average of the scores provided by these three LLM judges. To ensure consistency and reliability, we develop a comprehensive scoring guideline, which outline specific criteria for evaluating each component of the logging statement. These criteria address syntactic accuracy, semantic relevance, and contextual appropriateness, mitigating the limitations of traditional metrics like BLEU-4 and ROUGE-L.
Score Guideline 0: (Unacceptable) The logging statement is syntactically incorrect or misplaced. Formatting deviates significantly, and code changes may impair functionality or maintainability. 1: (Significant Issues) The statement is syntactically correct and appropriately placed but has major flaws: vague static text, incorrect log level, or missing key variables. Formatting inconsistencies or minor code alterations reduce readability but preserve functionality. 2: (Mostly Correct with Minor Flaws) The statement is semantically accurate, includes most relevant details, and uses an appropriate log level. Minor issues include verbose text, suboptimal formatting, or slight stylistic deviations. Functionality is preserved with trivial changes. 3: (Highly Accurate) The statement is precise, concise, and matches the ground truth. It uses the correct log level, parameterized logging, and consistent styling. The code retains full functionality and maintainability.
3.6. Implimentation Details
To evaluate conventional logging approaches and LLMs, we reproduced conventional methods using replication packages provided by their authors. For LLMs, we generated logging statements by calling their official APIs, setting the temperature to 0 to ensure deterministic outputs for identical queries, thus guaranteeing reproducibility. For SOLMs, we performed fine-tuning for one epoch on a dedicated dataset, using a batch size of 64 and a learning rate of 1e-4. All experiments, including training, fine-tuning, and inference, were conducted on a single NVIDIA A100 80GB GPU provided by Modal (Modal Labs, Inc., 2024), a serverless cloud infrastructure platform that supports easy deployment and reproducibility through provided scripts. Detailed experimental settings are available in our replication package (Renyi Zhong, Yichen Li, Guangba Yu, Wenwei Gu, Jinxi Kuang, Yintong Huo, Michael R. Lyu, 2025).
4. Empircial Study Results
4.1. RQ1: What are the most effective prompting strategies for using SOLMs in automated logging?
Motivation. Initially, our objective is to determine if the manner in which we engage with the model during the inference phase can have a profound effect on its success in generating automated logging statements. The structuring of the input prompt plays a crucial role in shaping how the SOLM interprets the given task and how it subsequently produces its outputs.
Approach. In addressing this question, our study is focused on evaluating the effectiveness of current prompting techniques within the domain of log generation. We specifically analyze, based on the definitions laid out in Section 3.3, the effectiveness of several prompting strategies: basic instruction prompting (base), in-context learning (ICL), retrieval-augmented generation (RAG), and chain-of-thought (CoT). The specific details regarding the prompt templates used can be found in Figure 3. To strengthen the generalizability of our findings, we employ multiple 7B instruction-following models, namely LLaMA3, Mistral, CodeLlama, and Qwen2.5-Coder, all in their original configurations.
Figure 3. The prompt template for automated logging.
Results. The quantitative evaluation comparing the four prompting techniques across the four selected 7B SOLMs is presented in Table 2. Our analysis of these results yields two primary findings concerning the performance of un-fine-tuned models and the efficacy of different prompt strategies for automated logging generation.
First, we observe a distinct performance disparity between the general-purpose instruction-following models (LLaMA3, Mistral) and the code-specific models (CodeLlama, Qwen2.5-coder) when used without any task-specific fine-tuning, i.e., the general-purpose models generally demonstrate superior performance on logging. For instance, LLaMA3 and Mistral achieve peak PA scores of 21.01 and 18.15, respectively (both using RAG), substantially higher than the peak PA scores achieved by CodeLlama (11.62 with ICL) and Qwen2.5-coder (13.25 with CoT). We attribute this trend to the weaker instruction-following capabilities in the original code-specific models. Furthermore, upon analyzing the failure cases, we noted a higher tendency for CodeLlama and Qwen2.5-coder to refuse the prompt or return empty responses compared to LLaMA3 and Mistral, which negatively impacts their effectiveness in this experimental setting. This suggests that, without fine-tuning, the broader instruction comprehension of general-purpose models may be more advantageous for automated logging tasks.
Findings 1: Without fine-tuning, general-purpose models outperform code-specific models for automated logging due to their better instruction-following ability.
Second, RAG emerges as the most effective prompting technique for enhancing automated logging generation genertaion performance. While other techniques occasionally yield the top score for an isolated metric, RAG demonstrates the most significant and robust improvements across the majority of metrics and models. Notably, for Mistral, RAG achieves the highest scores across all evaluated metrics, including PA (18.15), F1 (40.21), and ROUGE-L (34.74). For LLaMA3, RAG secures the top performance in PA (21.01), PMR (41.68), F1 (46.28), BLEU-4 (15.95), and ROUGE-L (36.73). For CodeLlama and Qwen2.5-coder, RAG generally leads to substantial gains over the baseline, particularly in metrics like F1 (CodeLlama: 39.01, Qwen2.5-coder: 49.28) and ROUGE-L (CodeLlama: 34.36, Qwen2.5-coder: 37.57). These results strongly indicate that providing relevant contextual information retrieved from a knowledge base significantly aids the SOLMs in accurately determining where to log, what variables to include, and formulating appropriate log messages, making RAG a highly promising strategy.
Findings 2: RAG proves the most effective prompting technique, significantly enhancing automoated logging statement generation performance across models and metrics.
Table 2. Comparison of prompting techniques for automated logging generation using four 7B instruction-following SOLMs.
Model Techs Location Level Variable Text
PA LA AOD PMR F1 BLEU-4 ROUGE-L
LLaMA3 base 13.32 54.50 83.06 38.00 30.41 9.68 28.92
ICL 9.99 50.67 81.25 30.00 30.95 11.31 30.60
RAG 21.01 53.72 82.09 41.68 46.28 15.95 36.73
COT 6.16 28.65 72.64 10.27 19.15 6.96 23.49
Mistral base 12.85 46.89 79.86 25.13 23.70 9.91 27.19
ICL 11.99 40.56 77.08 31.11 28.78 8.05 26.26
RAG 18.15 65.14 86.35 37.25 40.21 14.36 34.74
COT 6.63 50.25 79.23 13.57 21.02 11.01 30.35
CodeLLAMA base 6.99 57.14 84.29 37.62 35.74 12.21 31.71
ICL 11.62 48.14 79.87 26.36 29.47 9.50 27.54
RAG 8.16 62.04 86.43 31.02 39.01 13.82 34.36
COT 5.06 47.37 82.09 20.39 35.63 11.44 29.49
Qwen2.5-coder base 4.30 58.91 83.14 27.91 29.55 8.61 27.13
ICL 2.03 45.90 78.69 36.07 43.50 8.64 31.21
RAG 3.40 59.80 83.33 45.10 49.28 15.10 37.57
COT 13.25 51.01 80.21 20.35 48.66 10.58 29.52
4.2. RQ2: What’s the best tuning strategy using SOLMs for automated logging?
Motivation. In the course of the fine-tuning operation, a diverse array of strategies has the potential to influence the efficacy of our tasks significantly. In order to systematically assess the capabilities of the SOLMs, we thoroughly investigate the following factors:
(RQ2.1) Which PEFT technique yields the optimal performance for automoated logging statement generation using SOLMs? While SOLMs are more compact than larger LLMs, fully fine-tuning them for specific downstream tasks like automated logging can still be computationally demanding and may risk overfitting on task-specific data. PEFT methodologies have emerged as a compelling solution, enabling adaptation by updating only a small fraction of the model’s parameters or by adding a small set of new, trainable parameters. This significantly reduces computational costs and avoids catastrophic forgetting of the model’s pre-trained knowledge. However, a diverse range of PEFT techniques exists, each employing different mechanisms to inject task-specific information into the SOLM. For automated logging, which involves understanding code context, identifying appropriate logging locations, and generating relevant log messages, it is unclear which PEFT strategy offers the optimal balance of performance and efficiency. Therefore, we conduct an evaluation of the performance of SOLMs fine-tuned with various prominent PEFT techniques to ascertain which techniques yield superior performance outcomes.
(RQ2.2) How does the size of SOLMs impact the performance-resource trade-offs in automated logging? Although we focus on SOLMs, there is still considerable variation in size within this class. Larger models might capture more complex code patterns and lead to higher accuracy, but they could also incur greater computational costs during fine-tuning and inference. Therefore, evaluating the impact of model size is essential to understand the performance-resource trade-offs specific to automated logging.
(RQ2.3) Does the instruct variant of a SOLM outperform its base counterpart for automated logging? The distinction between using a ‘base’ pre-trained model versus an ‘instruct’ version of an SOLM presents another critical strategic choice. Instruct models are tuned to follow user prompts and formats, which is useful for specific tasks, but their tuning is generally broad. Base models, on the other hand, represent the raw capabilities learned during pre-training and might offer greater plasticity when fine-tuned on a highly specific downstream task like ours. Therefore, we aim to clarify which model mode serves as a better foundation for this task.
Approach. To address RQ2.1, we selected the 7B parameter versions of all four SOLMs as our primary subjects for investigating the impact of different PEFT techniques. We systematically evaluated four prominent PEFT methods: Prefix Tuning, Prompt Tuning, LoRA, and QLoRA. To establish a comparative baseline, we also measured the performance using direct inference without any PEFT fine-tuning (referred to as ‘base’). For all experiments conducted under RQ2.1, including the baseline, we consistently utilized the RAG-enhanced prompt format that has been identified as effective during our experiment in RQ1.
For RQ2.2, focusing on the influence of model size, we selected the Qwen2.5-coder model series. This choice is driven by the public availability of multiple versions within the same model family, specifically those with 0.5B, 1.5B, 3B, 7B, and 14B parameters, enabling a controlled comparison. Based on the findings from RQ2.1 where LoRA demonstrated superior performance among the PEFT techniques, we exclusively employ LoRA for fine-tuning across these different model sizes. Furthermore, we evaluat each size using both the basic prompt (‘base’) and the RAG-enhanced prompt (‘RAG’), in order to explore whether the effectiveness of RAG varies with model scale, particularly to assess the RAG capabilities of the smaller SOLMs.
To address RQ2.3, we compare the performance of ‘base’ models against their ‘instruct’ counterparts for automated logging. We select the Mistral 7B model, available in both base and instruct variants, for a controlled comparison, as it got the best perfomance in RQ2.1. Both model modes are fine-tuned using LoRA and employed the RAG-enhanced prompt, consistent with the result of RQ2.1 and RQ2.2. The fine-tuning process spanned five epochs, and performance are evaluated using metrics including Reject Rate, PA, LA, AOD, PMR, F1, BLEU-4, and ROUGE-L. The Reject Rate, defined as the percentage of test set examples rejected by the fine-tuned model during inference due to misalignment with learned task-specific criteria, quantifies the model’s selectivity, reflecting its ability to adhere to prompt instructions and generate valid logging statements.
Table 3. Performance Comparison of PEFT Techniques for four 7B SOLMs in automated logging.
Model PEFT Techs Location Level Variable Text
PA LA AOD PMR F1 BLEU-4 ROUGE-L
LLaMA3 base 21.01 53.72 82.09 41.68 46.28 15.95 36.73
prefix tuning 20.98 57.46 83.59 37.94 45.27 14.42 35.71
prompt tuning 25.01 67.38 87.85 43.01 50.33 15.66 35.96
LoRA 56.84 63.56 87.37 50.15 58.22 19.84 40.89
QLoRA 45.27 57.18 83.56 44.66 51.29 15.97 36.60
Mistral base 18.15 65.14 86.35 37.25 40.21 14.36 34.74
prefix tuning 20.35 63.34 85.20 38.95 43.07 15.96 36.75
prompt tuning 31.87 60.92 85.41 43.68 53.59 17.81 38.36
LoRA 63.97 69.50 88.64 52.79 57.89 23.40 45.73
QLoRA 61.90 69.28 88.66 53.31 58.97 22.42 44.70
CodeLlama base 8.16 62.04 86.43 31.02 39.01 13.82 34.36
prefix tuning 16.35 62.12 86.07 41.14 42.43 17.19 38.48
prompt tuning 18.15 62.42 86.81 41.76 49.91 18.41 39.69
LoRA 59.01 68.57 88.74 52.77 59.10 23.10 44.92
QLoRA 58.97 68.27 88.62 52.00 58.12 22.30 44.48
Qwen2.5Coder base 3.40 59.80 83.33 45.10 49.28 15.10 37.57
prefix tuning 25.27 65.48 87.80 44.14 56.36 15.23 35.47
prompt tuning 30.04 67.63 88.01 46.12 54.63 18.80 38.74
LoRA 62.40 68.46 88.82 52.24 58.98 23.04 44.96
QLoRA 60.21 67.87 88.70 51.55 59.47 22.52 44.03
Results. (RQ2.1) Table 3 shows the performance comparison of PEFT techniques for SOLMs, where we observe that all evaluated PEFT methods consistently improve performance over the baseline across all SOLMs and most metrics. For instance, looking at the prediction accuracy, QLoRA fine-tuning increased PA from 13.32 to 45.27 for LLaMA3, from 12.85 to 61.90 for Mistral, from 6.99 to 58.97 for CodeLlama, and from 4.30 to 60.21 for Qwen2.5Coder. Similar substantial gains are observed across other metrics like F1 score for variable prediction and BLEU-4/ROUGE-L for statement generation, indicating that fine-tuning with parameter-efficient techniques, is crucial for adapting SOLMs to this specific task.
Furthermore, LoRA emerges as the most effective PEFT methodology, achieving the highest scores for the majority of metrics across all four models. For example, with LoRA, LLaMA3 gets the top PA (56.84), PMR (50.15), F1 (58.22), BLEU-4 (19.84), and ROUGE-L (40.89). Similarly, LoRA leads to the best PA (63.97), LA (69.50), BLEU-4 (23.40), and ROUGE-L (45.73) for Mistral. CodeLlama shows LoRA as the top performer across all metrics. For Qwen2.5Coder, LoRA is also dominant, securing the best PA (62.40), LA (68.46), AOD (88.82), PMR (52.24), BLEU-4 (23.04), and ROUGE-L (44.96). Moreover, QLoRA is competitive to LoRA, achieving the second-best results or even surpassing LoRA in a few specific instances (e.g., F1 for Qwen2.5Coder, PMR, and F1 for Mistral). Prompt Tuning shows some strength, particularly for ‘level’ prediction with LLaMA3, while Prefix Tuning, though an improvement over the baseline, is generally outperformed by LoRA, QLoRA, and Prompt Tuning.
Additionally, the venn diagram in Figure 4(a) illustrates the overlap of corrected logging statement placements across different PEFT variations for Mistral-7B. The largest overlap (183) is observed in the central region, indicating a core set of logging statements consistently corrected across all PEFT methods (base, Prefix Tuning, Prompt Tuning, LoRA, and QLoRA). LoRA and QLoRA show significant individual contributions (125 and 86, respectively), suggesting their effectiveness in identifying unique logging placements, while the base method contributes the least (40), highlighting the improvement brought by PEFT techniques.
Findings 3: Fine-tuning with PEFT techniques significantly enhances SOLMs performance for automated logging, with LoRA demonstrating the most consistent and superior results across the evaluated models and metrics.
(a) Mistral-7B with PEFT Variations
(b) Qwen2.5-coder of Different Sizes
(c) Mistral-Instruct-7B Across Epochs
Figure 4. Overlap of Corrected Logging Statement Placement Across Different SOLM Configurations.
Table 4. Impact of Model Size on Performance and Resource Usage for automated logging.
model params trainable params training time (s/epoch) inference time (s/prompt) PA LA AOD PMR F1 BLEU-4 ROUGE-L
0.5B ~281M 2438.2749 0.0959 21.38 61.53 86.78 41.74 56.59 18.95 39.55
1.5B ~485M 4395.2893 0.1866 53.11 60.69 85.70 50.34 58.39 20.35 40.69
3B ~652M 7694.3684 0.2375 52.18 68.41 88.77 49.20 56.94 22.23 43.18
7B ~1139M 13258.3642 0.2710 62.40 68.46 88.82 52.24 58.98 23.04 44.96
14B ~1625M 21780.5427 0.4854 66.20 69.92 89.36 54.53 59.93 25.20 47.22
(RQ2.2) Table 4 reveals that increasing model size leads to improved performance in automated logging, particularly for models 3B and larger. Across almost all metrics, there is a discernible improvement as the model parameter count increases from 0.5B to 14B. For example, prediction accuracy (PA) improves from 21.38 (0.5B) to 66.20 (14B), and ROUGE-L scores for text generation increase from 39.55 (0.5B) to 47.22 (14B). The 14B model consistently outperforms all smaller variants, and the 7B model also shows strong performance.
Models smaller than 3B (i.e., 0.5B and 1.5B) exhibit less stable performance scaling. While the 0.5B model is generally the weakest, the progression to the 1.5B and then to the 3B model is not uniformly positive across all metrics. For instance, the 1.5B model shows a slight decrease in LA (60.69 vs 61.53 for 0.5B) and AOD (85.70 vs 86.78 for 0.5B). Furthermore, when moving from 1.5B to 3B, there are slight dips in PA (52.18 vs 53.11), PMR (49.20 vs 50.34), and F1 (56.94 vs 58.39). This suggests that while larger models generally perform better, the performance gains for models below 3B parameters might be less consistent or predictable for this specific task and fine-tuning approach.
Larger models come with increased resource requirements. As model size increases, the training time per epoch, and inference time per prompt also escalate substantially. For instance, training time per epoch rises from approximately 2438 seconds for the 0.5B model to 21780 seconds for the 14B model. Similarly, inference time per prompt increases from 0.0959 seconds (0.5B) to 0.4854 seconds (14B).
The venn diagram in Figure 4(b) depicts the overlap of corrected logging statement placements across different sizes of the Qwen2.5Coder model (0.5B, 1.5B, 3B, 7B, 14B). The central overlap (442) represents logging statements consistently corrected across all sizes, with the 14B model contributing the most unique placements (152), followed by 7B (73). Smaller models (0.5B and 1.5B) show limited unique contributions (21 and 6, respectively), reinforcing the finding that models with 3B+ parameters perform better for this task.
Findings 4: For models with 3B+ parameters, performance in generating logging statements improves with more parameters, but increases computational costs, indicating a performance-resource trade-off. Models under 3B show inconsistent scaling, suggesting a minimum capacity may needed for this task.
(a) Base vs. Instruct on Reject Rate
(b) Base vs. Instruct on PA
(c) Base vs. Instruct on PMR
(d) Base vs. Instruct on ROUGE-L
Figure 5. Performance Comparison of Base and Instruct Mistral-7B Models Across Five Epochs.
(RQ2.3) Figure 5 illustrates the performance comparison between the base and instruct variants of the Mistral-7B model across five epochs for automated logging. The instruct model consistently outperforms the base model across most metrics, including Reject Rate, PA, and ROUGE-L. For instance, the instruct model achieves a lower Reject Rate, indicating better adherence to task-specific criteria and fewer invalid outputs (Figure 5(a)). Similarly, the instruct model demonstrates higher PA (Figure 5(b)), reflecting superior accuracy in predicting logging statement placement. The ROUGE-L scores (Figure 5(d)) further confirm that the instruct model generates logging statements with greater textual similarity to the ground truth. These trends are evident from the first epoch and persist through the fifth, suggesting that the instruct model’s pre-training for instruction-following enhances its ability to adapt to our logging task.
Findings 5: The instruct variant of SOLM model outperforms its base counterpart in automated logging, benefiting from its instruction-tuned foundation, which enhances task adherence and output quality.
In addition, the performance trends across epochs reveal that excessive fine-tuning can lead to diminishing returns. For both models, the Reject Rate and PA peak at the first epoch, where the Reject Rate is minimized, and the number of correctly predicted logging statement placements is maximized (Figure 5(a) and Figure 5(b)). Beyond the first epoch, both metrics show a slight decline or stabilization, with the Reject Rate marginally increasing and PA slightly decreasing by the fifth epoch. This trend suggests that additional fine-tuning may lead to overfitting, causing the models to become overly specialized to the training data and potentially losing some generalizability. For the instruct model, this could also imply a partial erosion of its pre-trained instruction-following robustness.
The venn diagram in Figure 4(c) further illustrates the overlap of corrected logging statement placements across epochs for the Mistral-7B-Instruct model. The central overlap (1120) indicates a stable core of corrections maintained across all five epochs, with the first epoch contributing the most unique corrected placements (40). The decline in unique contributions from later epochs (e.g., 3 for the fifth epoch) supports the observation of diminishing returns and potential overfitting with prolonged fine-tuning.
Findings 6: Both base and instruct models achieve optimal performance at the first epoch for Reject Rate and PA, with prolonged fine-tuning leading to slight performance declines due to overfitting, indicating that excessive fine-tuning may compromise pre-trained capabilities.
4.3. RQ3: How effectively do SOLMs compare to existing methods and LLM baselines in automated logging?
Motivation. Having established optimal strategies for employing SOLMs in addressing the preceding RQs, this study aims to investigate the performance of SOLMs in automated logging compared to existing methods and the direct application of LLMs.
  • (RQ3.1) How effectively do these methods determine appropriate logging locations?
  • (RQ3.2) What is the quality of the logging statements generated by these methods?
  • (RQ3.3) When evaluated by an LLM acting as a judge, how does the overall quality of the logging statements produced by these methods compare?
Approach. To address RQ3, we evaluate the performance of SOLMs in automated logging against existing method (i.e., LANCE, LEONID, Unilog, and Fastlog) and LLMs (i.e., Claude3.7sonnet, Deepseek-coder-v3, GPT4o, and LLaMA3.1-405B). We assess the loging location accuracy (PA), the statement quality (LA, AOD, PMR, F1, BLEU-4, and ROUGE-L), and overall quality via our LLM judges. Target SOLMs (LLaMA-8B, Mistral-7B, CodeLlama-13B, Qwen2.5-coder-14B) are fine-tuned with LoRA and RAG, while LLMs were tested in base, ICL, RAG, and COT configurations. To ensure a comprehensive evaluation of SOLMs’ capabilities, we select the largest parameter models available within the SOLM definition (i.e., open-source models with fewer than 14B parameters). This choice maximizes the potential performance of SOLMs, allowing us to showcase their optimal effectiveness in automated logging.
Result. (RQ3.1) Table 5 shows the performance comparison of automated logging approaches across all metrics. The table shows that Qwen2.5-coder-14B-RAG-LoRA achieves the highest PA at 66.20%, outperforming all other models, including LLMs like Claude3.7sonnet-RAG (65.90%), Deepseek-coder-v3-RAG (65.20%), as well as existing methods like Fastlog (53.40%). Among SOLMs, Mistral-7B-RAG-LoRA (63.97%) and CodeLlama-13B-RAG-LoRA (63.57%) also surpass most LLMs and all existing methods, indicating that fine-tuning with LoRA and RAG significantly enhances the ability of SOLMs to identify logging locations in various configuration settings.
(RQ3.2) For statement quality, Qwen2.5-coder-14B-RAG-LoRA again leads across all metrics. These results surpass the best-performing LLM, deepseek-coder-v3-RAG (e.g., 68.39% LA, 18.97% BLEU-4), and existing methods like Fastlog (e.g., 59.26% LA, 13.28% BLEU-4). CodeLlama-13B-RAG-LoRA also performs strongly, particularly in BLEU-4 (24.39%) and ROUGE-L (46.91%), closely rivaling Qwen2.5-coder-14B. The superior performance of fine-tuned SOLMs suggests that targeted optimization enables them to generate more accurate, relevant, and contextually appropriate logging statements than both unoptimized LLMs and traditional methods.
(RQ3.3) Figure 6 illustrates the score distribution for LLM judges evaluating the quality of logging statements from various methods. First, the bar chart indicates the number of cases receiving each score (0, 1, 2, 3) for different models. Qwen2.5-coder-14B stands out with the highest number of cases scoring 3, alongside the lowest number of cases scoring 0. This distribution suggests that Qwen2.5-coder-14B consistently generates logging statements of higher quality. Second, the trend line representing the average score highlights Qwen2.5-coder-14B achieving the highest average score of 1.506, surpassing all other models, including Claude3.7sonnet-RAG (1.489) and Deepseek-coder-v3-RAG (1.467).
Figure 7 presents a case study comparing logging statements generated by different models for an error logging scenario in a Java method handling a DuplicateKeyException. The ground truth logging statement is log.error(”Update alert group error, groupName:{}”), alertGroup.getGroupName(), ex). Among the LLMs, Claude3.7sonnet-RAG generates a statement that closely resembles the ground truth but includes an additional variable (id), which may add unnecessary verbosity. Deepseek-coder-v3-RAG and GPT4o-RAG produce statements with missing or misused variables (e.g., using groupName directly instead of alertGroup.getGroupName()), reducing their contextual accuracy. LLaMA3.1-405B-RAG omits the exception variable (ex), limiting its diagnostic utility. In contrast, fine-tuned SOLMs demonstrate superior performance: Mistral-7B-RAG-LoRA and CodeLlama-13B-RAG-LoRA generate statements identical to the ground truth, ensuring both accuracy and relevance. Qwen2.5-coder-14B-RAG-LoRA adds minor additional text (”already exist”) but retains all critical components, closely aligning with the ground truth.
Table 5. Performance Comparison of Automated Logging Approaches Across All Metrics.
Location Level Variable Text
Model PA LA LOD PMR F1 BLEU-4 ROUGE-L
Existing Approach
LANCE 44.67 48.23 80.60 26.84 48.33 11.21 29.38
LEONID 46.74 49.67 81.88 28.04 49.65 12.63 31.33
Unilog 51.27 56.34 84.58 34.19 50.46 14.05 34.68
Fastlog 53.40 59.26 85.23 38.22 51.24 13.28 32.54
Large Language Model
Claude3.7sonnet-base 47.02 66.22 87.83 46.25 55.14 17.32 39.85
Claude3.7sonnet-ICL 62.80 62.46 86.64 44.70 55.06 14.93 36.84
Claude3.7sonnet-RAG 65.90 65.89 88.01 45.78 56.49 17.39 39.64
Claude3.7sonnet-COT 47.32 64.60 87.95 35.33 55.58 15.13 36.87
Deepseek-coder-v3-base 53.65 67.85 88.04 48.23 55.14 17.42 38.40
Deepseek-coder-v3-ICL 62.64 64.06 86.89 46.94 53.76 15.98 37.07
Deepseek-coder-v3-RAG 65.20 68.39 88.08 49.28 55.93 18.97 41.32
Deepseek-coder-v3-COT 49.05 65.11 87.31 34.96 56.12 14.31 35.32
GPT4o-base 27.91 63.60 86.89 47.73 52.62 15.66 36.95
GPT4o-ICL 52.45 56.44 84.63 44.51 52.96 13.02 32.56
GPT4o-RAG 55.78 63.28 86.75 46.27 54.20 15.66 36.94
GPT4o-COT 24.84 62.73 86.91 38.47 56.22 15.03 36.46
LLaMA3.1-405B-base 46.45 63.80 86.96 51.76 54.70 18.50 39.37
LLaMA3.1-405B-ICL 51.62 55.10 83.25 46.45 52.01 15.01 35.71
LLaMA3.1-405B-RAG 55.04 63.64 86.94 51.06 56.04 19.55 40.94
LLaMA3.1-405B-COT 15.25 60.92 85.23 39.74 50.94 18.22 37.71
Fine-tuned Small Open-source Language Models
LLaMA3-8B-RAG-LoRA 56.84 63.56 87.37 50.15 58.22 19.37 42.03
Mistral-7B-RAG-LoRA 63.97 69.50 88.64 52.79 57.89 20.62 41.70
CodeLlama-13B-RAG-LoRA 63.57 64.43 87.17 53.90 58.59 24.39 46.91
Qwen2.5-coder-14B-RAG-LoRA 66.20 69.92 89.36 54.53 59.93 24.05 46.51
Figure 6. Distribution of LLM Judges Scores for Generated Logging Statement Quality Across Models.
Figure 7. A Case of Generated Logging Statement from Multiple Models.
Findings 7: Fine-tuned SOLMs outperform both existing methods and LLMs across all evaluated metrics, demonstrating superior logging location accuracy and statement quality.
4.4. RQ4: Can SOLMs generalize logging statement generation across diverse code repositories?
Motivation. In real-world software development, logging practices vary significantly across projects due to differences in coding styles, project domains, and developer preferences. For SOLMs to be practically viable for automated logging, they must demonstrate robust generalization when applied to unseen repositories. Poor generalization could lead to ineffective logging statements that fail to capture critical runtime information or adhere to project-specific conventions, limiting the utility of SOLMs in cross-project settings.
Approach. To evaluate the cross-repository generalization ability of SOLMs, we design an experiment that trains models on a subset of repositories from the dataset and tests them on a distinct, non-overlapping set of repositories. We randomly partition the dataset into two groups, each containing five repositories, ensuring diversity in project domains. In each experimental run, we use one group of five repositories as the training set, three repositories from the other group as the validation set, and the remaining two repositories from the same group as the test set. The validation set primarily served to support RAG by providing a pool of examples from which the most similar code snippets are retrieved using the BM25 algorithm. This partitioning ensures that the test set represents unseen projects with potentially different coding styles and logging conventions, simulating real-world cross-repository scenarios. The experimental setup involves training on a subset of repositories (R1, R3, R5, R7, R9). It also includes validating on another subset (R2, R4, R6) and testing on unseen repositories (R8, R10) in the first configuration. The second configuration, meanwhile, trains on Apache-dominated repositories (R1, R3, R5, R7, R9) and tests on non-Apache repositories. For this experiment, we select the two best-performing 7B SOLMs from RQ2.1: Mistral and Qwen2.5-coder. We fine-tune these models using the best-performing strategy identified in prior experiments, combining LoRA with RAG.
Results. Table 6 presents the performance of two fine-tuned SOLMs, Mistral-7B and Qwen2.5-coder-7B, in automated logging across diverse code repositories. The results demonstrate the generalization capabilities of SOLMs and highlight the impact of similar logging practices on cross-project performance.
Table 6. Generalization Capabilities for SOLMs Across Diverse Code Repositories.
Train Data Valid Data Test Data Model PA LA AOD PMR F1 BLEU-4 ROUGE-L
R1, R3, R5, R7, R9 R2, R4, R6 R8, R10 Mistral 52.97 69.30 90.74 44.31 55.87 18.80 40.15
Qwen2.5coder 55.54 71.67 91.86 44.67 57.20 20.21 42.22
R2, R4, R6, R8, R10 R1, R3, R5 R7, R9 Mistral 44.27 51.76 79.75 43.44 47.18 17.25 38.46
Qwen2.5coder 50.71 56.19 81.96 45.71 53.30 18.48 40.50
The result shows that both Mistral-7B and Qwen2.5-coder-7B exhibit robust performance. Specifically, Qwen2.5-coder achieves a PA of 55.54% and ROUGE-L of 42.22%, while Mistral-7B achieves comparable results with a PA of 52.97% and ROUGE-L of 40.15%. These metrics indicate that both models successfully generate accurate logging statements and identify appropriate logging locations in unseen repositories, even when the test set includes projects with distinct logging conventions. The high AOD (91.86% for Qwen2.5-coder) and ROUGE-L scores suggest that the generated logs closely align with ground-truth statements in terms of log level and text similarity. This reveals that SOLMs, when fine-tuned with LoRA and RAG, proficiently generalize across various project areas without experiencing a notable decline in performance.
Findings 8: SOLMs demonstrate strong generalization capabilities in automated logging, maintaining high performance across diverse, unseen repositories.
The performance difference between the two configurations in Table 6 highlights the influence of similar logging practices on cross-project generalization. In the first configuration, where both training and test sets include Apache open-source projects, the models achieve significantly higher performance (e.g., Qwen2.5-coder: PA 55.54%, ROUGE-L 42.22%) compared to the second configuration, where the training set comprises Apache projects, but the test set includes non-Apache projects (e.g., Qwen2.5-coder: PA 50.71%, ROUGE-L 40.50%). The performance drop in the second configuration (e.g., 4.83% lower PA and 1.72% lower ROUGE-L for Qwen2.5-coder) suggests that the absence of shared logging conventions, such as those prevalent in Apache projects (e.g., consistent verbosity levels and formatting styles), reduces the models’ ability to generate contextually appropriate logs. This may because Apache projects follow standardized logging guidelines, aiding knowledge transfer during fine-tuning, while non-Apache projects may use diverse, project-specific practices that hinder generalization. This disparity underscores that similarity in logging practices between repositories can enhance cross-project performance.
Findings 9: Fine-tuning SOLMs with data reflecting similar logging practices, such as those prevalent in Apache open-source projects, significantly enhances their cross-project generalization capabilities.
5. Discussion
Based on previous experiments, this section analyzes the strengths of using SOLMs and elicits potential future research directions to advance logging practice.
5.1. Analysis of SOLMs’ Advantages
5.1.1. Efficiency and Cost-Effectiveness
SOLMs exhibit remarkable efficiency and cost-effectiveness in automated logging, making them a compelling alternative to LLMs. Unlike LLMs, which often require extensive computational resources, including thousands of GPU hours for training and inference, SOLMs achieve comparable performance with significantly lower resource demands. For instance, our experiments demonstrate that the Qwen2.5-coder-14B model, with 14 billion parameters, can be fine-tuned within 6 hours using a single Nvidia A100 GPU, producing high-quality logging statements with a PA of 66.20% and a ROUGE-L score of 46.51% (Table 5). This efficiency reduces hardware costs and energy consumption, aligning with sustainable computing goals.
5.1.2. Privacy and Security
Privacy preservation is a standout advantage of SOLMs. Li et al. (Li et al., 2024a) highlight that LLMs often rely on cloud-based APIs, posing risks of proprietary code leakage. In contrast, SOLMs’ smaller size enables local deployment, ensuring sensitive code remains secure. This is particularly valuable for companies where strict data protection is paramount, offering a safer alternative for logging generation.
5.1.3. Adaptability to Enterprise-specific styles
One of the key challenges in automated logging is adapting to enterprise-specific logging styles and conventions, which vary significantly across organizations due to differences in verbosity levels, error prioritization, or compliance-driven formatting. Our findings demonstrate that SOLMs, when fine-tuned with techniques such as LoRA and RAG, can effectively align with project-specific logging practices. This adaptability is particularly valuable in real-world scenarios where organizations maintain proprietary logging guidelines. Unlike general-purpose LLMs, which struggle to adapt without extensive retraining, SOLMs can be fine-tuned efficiently on internal codebases, ensuring alignment with organizational standards while minimizing computational overhead.
5.2. Future Work Directions
5.2.1. Integration into Development Tools
To maximize the practical impact of SOLMs in automated logging, future work should focus on integrating these models into widely used development tools, such as integrated development environments (IDEs), and CI/CD platforms. Real-time logging statement suggestions during code authoring or automated insertion during code reviews could streamline the development process and reduce manual effort. For instance, an IDE plugin leveraging a fine-tuned SOLM could analyze code context on-the-fly, recommend logging points, and suggest high-quality logging statements tailored to the project’s conventions. Such integrations would require optimizing SOLMs for low-latency inference and ensuring compatibility with diverse development workflows. Additionally, incorporating user feedback mechanisms into these tools could enable iterative refinement of generated logs, further aligning them with developer preferences.
5.2.2. Addressing Dynamic Logging Requirements
Logging practices often evolve during a project’s lifecycle due to changing requirements, such as new debugging needs or compliance regulations. SOLMs must be capable of adapting to these dynamic requirements without requiring extensive retraining. Future research could investigate continual learning techniques to enable SOLMs to incrementally adapt to new logging conventions or project-specific requirements. For instance, online fine-tuning approaches could allow SOLMs to learn from newly added logging statements in a repository, ensuring sustained alignment with evolving practices. Additionally, exploring active learning strategies, where SOLMs query developers for feedback on ambiguous logging scenarios, could further enhance their adaptability.
6. Threats to Validity
6.1. Internal Validity
Selection of hyperparameters for fine-tuning. A potential threat to internal validity lies in the selection of hyperparameters for fine-tuning the SOLMs (e.g., learning rate, batch size, prefix length for prefix tuning, rank for LoRA). The study utilized recommended hyperparameters from official sources due to their proven effectiveness. However, these hyperparameters may not be optimal for all models or datasets, potentially introducing bias in the performance results. Suboptimal hyperparameter choices could lead to underperformance or overfitting, affecting the observed effectiveness of SOLMs compared to baseline LLMs or existing methods. To mitigate this, we conducted preliminary experiments to validate the chosen hyperparameters on a subset of the AL-Bench dataset, ensuring reasonable performance. Nonetheless, a more exhaustive hyperparameter search may further enhance model performance and reduce this threat. The potential data leakage. A potential threat to the internal validity of our study is the possibility that the AL-Bench dataset, which comprises logging statements from 10 widely used open-source projects, may have been included in the pre-training corpora of the SOLMs or LLMs evaluated. Since these models are pre-trained on large-scale datasets, often including publicly available code repositories from platforms like GitHub, there is a risk that some or all of the AL-Bench projects were part of their training data. Such data leakage could artificially inflate the performance of these models, particularly in zero-shot or few-shot settings, potentially skewing our results and conclusions. To mitigate this threat, we carefully analyzed the performance of base (non-fine-tuned) models in our experiments. The results, as shown in Table 3, indicate that base models, such as Qwen2.5-coder (PA: 4.30%), exhibit significantly lower performance compared to their fine-tuned counterparts (PA: 62.40%). This substantial performance gap suggests that the base models, despite potential exposure to AL-Bench data during pre-training, do not inherently possess the task-specific knowledge required for high-quality automated logging. Instead, the superior performance of fine-tuned SOLMs is primarily attributable to our fine-tuning strategies (e.g., LoRA and RAG), which adapt the models to the specific logging task and dataset. Thus, we argue that any potential data leakage has minimal impact on the validity of our conclusions, as the observed improvements stem from task-specific fine-tuning rather than pre-existing knowledge from pre-training. Nonetheless, to further address this threat in future work, we recommend evaluating models on proprietary or newly created datasets that are guaranteed to be absent from pre-training corpora.
6.2. External Validity
The representativeness of the dataset. A potential threat to generalizability is that the AL-Bench dataset, comprising 10 open-source projects, may not fully represent logging practices in proprietary codebases. We mitigated this by selecting projects from diverse domains (e.g., task scheduling, messaging systems, IoT platforms), ensuring broad coverage of logging requirements. However, similar to prior work (Li et al., 2024a), our dataset is predominantly Java-based, which may limit the generalizability of SOLM’s performance in generating log statements for other programming languages. This language-specific focus could restrict insights into how SOLM performs across diverse language ecosystems, potentially affecting its applicability in non-Java contexts. The selection of SOLM. Another potential threat to the generalizability of our findings lies in the selection of specific SOLMs evaluated in this study. While these models were chosen based on their established performance in software engineering tasks, they may not fully represent the diversity of available SOLMs or future advancements in model architectures. For instance, other SOLMs with different pre-training datasets, architectural designs (e.g., transformer variants or mixture-of-experts models), or domain-specific optimizations might exhibit varying performance in automated logging. To mitigate this threat, we selected models with broad applicability in code-related tasks and ensured they were fine-tuned using techniques (e.g., LoRA and RAG) to align with logging-specific requirements. However, future work should explore a wider range of SOLMs, including those with different training corpora or specialized architectures, to validate the robustness of our findings across diverse model ecosystems.
6.3. Construct Validity
Adequacy of evaluation metrics for logging quality. A potential threat to construct validity lies in whether the chosen evaluation metrics fully capture the quality of generated logging statements. These metrics primarily assess syntactic similarity to ground-truth logs (e.g., ROUGE-L for text similarity and correctness of logging location (PA). However, high-quality logging statements must also provide actionable insights for developers, such as facilitating debugging or system monitoring, which may not be fully reflected by these metrics. For instance, a generated log might score high on ROUGE-L due to textual similarity but fail to capture critical runtime context (e.g., omitting key variables or using an inappropriate verbosity level). To mitigate this, we incorporated the LLM-as-a-judge approach to assess overall quality holistically. Nevertheless, future work could include developer-centric evaluations, such as user studies, to validate the practical utility of generated logs. Representativeness of the LLM judger result. The use of an LLM to evaluate the quality of generated logging statements introduces a construct validity threat if the LLM’s judgments do not align with human developer preferences. While the LLM-as-a-judge approach has shown promise in software engineering tasks (Wang et al., 2025), its scoring may not fully capture nuanced aspects of log quality, such as clarity, relevance to specific debugging scenarios, or adherence to project-specific logging conventions. Misalignment between LLM and human judgments could lead to over- or underestimation of SOLM performance. Therefore, incorporating human evaluations or domain-specific rubrics in future work could enhance the alignment between the LLM judger and practical logging needs.
7. Related Work
7.1. Automated Logging
Traditionally, the automation of logging statements is divided into two primary stages (Chen and Jiang, 2021; He et al., 2021): the identification of logging locations and the creation of logging statements. These stages are respectively denoted as where to log and what to log (Zhong et al., 2025). In addressing the complexities associated with determining where to log, various methodologies have been investigated by researchers to identify appropriate logging locations within source code (Jia et al., 2018; Li et al., 2020a; Yao et al., 2018; Zhao et al., 2017; Zhu et al., 2015; Yuan et al., 2012b; Li et al., 2018). Regarding what to log, the generation of logging statements is usually segmented into three specific subtasks: the generation of logging text (Ding et al., 2022), the selection of logging variables (Liu et al., 2019; Yuan et al., 2012c), and the prediction of the logging level (Li et al., 2017, 2021a; Liu et al., 2022; Mizouchi et al., 2019).
The latest methodology offers a solution for the automatic generation of logging statements, addressing the selection of logging locations, determining the levels of statements, composing content, and identifying variables in a single step. Mastropaolo et al. (Mastropaolo et al., 2022) introduced LANCE, a pioneering comprehensive tool that creates complete logging statements powered by T5. In addition to this, they developed LEONID (Mastropaolo et al., 2024), which integrates deep learning with information retrieval techniques to enhance performance. Meanwhile, Xu et al. (Xu et al., 2024) presented UniLog, grounded in the in-context learning framework of LLMs. Additionally, Xie et al. (Xie et al., 2024) introduced FastLog, which is capable of swiftly generating and inserting entire logging statements. Furthermore, Li et al. (Li et al., 2024b) proposed SCLogger, noted as the first approach to generate contextualized logging statements utilizing inter-method static context.
In this paper, we distinguish our work by focusing on the use of SOLMs for automated logging, addressing the limitations of LLMs in terms of privacy, computational efficiency, and adaptability to enterprise-specific logging practices. Unlike prior proposed approaches, which predominantly rely on LLMs, our study leverages fine-tuned SOLMs. This enables local deployment, mitigating privacy risks associated with cloud-based LLM APIs, and significantly reduces computational overhead, aligning with sustainable computing goals. Furthermore, our comprehensive evaluation using the AL-Bench dataset (Tan et al., [n. d.]) demonstrates SOLMs’ robust generalization across diverse, unseen repositories, a critical aspect not extensively explored in prior work. By systematically investigating prompting strategies and fine-tuning techniques, we provide a scalable and practical solution for automated logging that balances performance with resource constraints, offering a viable alternative for real-world software development.
7.2. Studies in Logging Practices
Advancements in logging within software engineering have sparked a growing interest in exploring logging practices across various domains. Zeng et al. (Zeng et al., 2019) and Chen(Chen and Jiang, 2017) extended the work of Yuan et al. (Yuan et al., 2012a) by analyzing log statements in Android and Java systems, revealing the widespread occurrence of logging in these environments. Kabinna et al. (Kabinna et al., 2016) investigated how changes such as bug fixes, feature enhancements, and code refactoring often lead to revisions in logging statements. Lai et al. (Lal et al., 2015) provided insights into logging code constructs at both file-level and block-level, addressing nine key research questions focused on statistical and content analysis. Li et al. (Li et al., 2021b) conducted a detailed qualitative study of the advantages and challenges associated with logging in software development, while Zhou et al. (Zhou et al., 2020) explored the connection between logging practices and data leaks in mobile applications. Zhao et al. (Zhao et al., 2023) analyzed IDs within log statements, proposing a straightforward approach to inject IDs to reduce information loss and examining the extent of information gained through this technique. Li et al.(Li et al., 2023b) investigated the characteristics and practical importance of dynamic variables, proposing a variable-aware log abstraction technique. Li et al.(Li et al., 2024a) introduced a study on LLM-assisted logging statement generation, demonstrating that prompt-based zero-shot or few-shot learning significantly enhances the generalization capabilities of LLMs. Tan et al. (Tan et al., [n. d.]) proposed AL-Bench, which is a comprehensive benchmark designed specifically for automatic logging tools.
8. Conclusion
In this paper, we explore solutions to the challenges associated with manual logging statement generation by comprehensively investigating the potential of SOLMs as a more viable alternative to resource-intensive and privacy-concerning LLMs. We conduct an extensive, large-scale empirical study to systematically evaluate the efficacy of four prominent SOLMs. This evaluation encompasses various prompt strategies, such as RAG; PEFT techniques, with a focus on LoRA; the impact of different model sizes; and the comparative performance of base versus instruction-tuned model types.
Our key findings provide several critical insights into leveraging SOLMs for automatic logging. We demonstrate that (1) RAG significantly enhances the performance of SOLMs in generating logging statements, and (2) LoRA proves to be a highly effective PEFT technique, enabling substantial improvements with minimal trainable parameters. (3) While larger SOLMs generally yield better results, this is balanced by increased computational demands, highlighting an important performance-resource trade-off. Furthermore, (4) instruction-tuned SOLMs consistently outperform their base counterparts, benefiting from their inherent instruction-following capabilities. Most notably, (5) our research establishes that fine-tuned SOLMs, particularly Qwen2.5-coder-14B, can surpass existing specialized logging tools and even larger LLM baselines in both the accuracy of predicting logging locations and the quality of the generated statements. These findings are further corroborated by LLM-as-a-judge evaluations, which confirm the high quality of SOLM-generated outputs. Additionally, (6) we find that SOLMs exhibit robust generalization capabilities across diverse, unseen code repositories, underscoring their practical applicability.
In conclusion, this study provides strong evidence that appropriately fine-tuned SOLMs offer a powerful, efficient, privacy-preserving, and adaptable solution for automated logging. By making our methodology, datasets, and results publicly available (Renyi Zhong, Yichen Li, Guangba Yu, Wenwei Gu, Jinxi Kuang, Yintong Huo, Michael R. Lyu, 2025), we aim to stimulate further research and development in this domain, ultimately contributing to improved software maintenance practices.
Acknowledgement
The work described in this paper was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14206921 of the General Research Fund).